OpenStudy (anonymous):

What is y=-2x^2-4x+3 in intercept form? Someone help, please?!

OpenStudy (anonymous):

you mean slope intercept form?

OpenStudy (anonymous):

Yes! @helpme1.2

OpenStudy (anonymous):

slope intercept form is y = mx+b, i don't think that function can be turned into slope form but you can turn it in to intercept form you could complete the square and turn in in to vertex form that could get it in to intercept form if i am not mistaken y=-2x^2-4x+3 y = -(2x^2+4x-3) y = -2(x^2+4x + 4)= 3+4 y = -2(x+2)^2 = 7 y = -2(x+2)^2 - 7= 0

OpenStudy (triciaal):

slope-intercept y = mx + b point-slope y-y1 = m(x-x1) standard form Ax + By = C